OpenStudy (anonymous):

compute the derivative of f(x) = (x^2 -3x- sqrt(x)/ sqrt(x)

OpenStudy (anonymous):

please show all calc involved

OpenStudy (rmrjr22):

F'(x)=-3+2 x

OpenStudy (rmrjr22):

x^2 = 2x , -3x = -3, sqrt(x)/sqrt(x) = 1, 1= 0

OpenStudy (bahrom7893):

i dont think she meant that

OpenStudy (bahrom7893):

i think she meant all of the top over all of the bottom

OpenStudy (anonymous):

u = x² -3x -√x --> u' = 2x -3 - 1/2√x v = √x ---> v' = 1/2√x = (2x -3 - 1/2√x) √x - [ (x² -3x -√x) 1/2√x ] / x = 3x (x -1)/ 2x^(3/2)

OpenStudy (rmrjr22):

f'(x) = (3 x-3)/(2 sqrt(x))

OpenStudy (anonymous):

Common denominator is 2√x

OpenStudy (anonymous):

Jinnie, I don't have that version of MS office on this laptop!

OpenStudy (mertsj):

\[\frac{x^2-3x-\sqrt{x}}{\sqrt{x}}\]

OpenStudy (anonymous):

@Jinne, could you check my u = ... , v = .... If it's right then my answer's correct!

OpenStudy (anonymous):

Thanks a bunch Mertsj :), then I'm correct!

OpenStudy (rmrjr22):

So my answer is correct, yes?

OpenStudy (anonymous):

I'm double check my work, rmrj!

OpenStudy (anonymous):

can anyone view the attachment I posted

OpenStudy (rmrjr22):

yes i saw ur document, couldnt see what your choice for D) was

OpenStudy (anonymous):

my choice for d and e are 2x-1, 2x-3, respectively

OpenStudy (anonymous):

I can't! Could you just type it up!

OpenStudy (anonymous):

cholorphyll. give me a bit of time. i can switch computers. i can't type it out because its too funky.

OpenStudy (anonymous):

I'm double check my calculation, anyway! I just broke my powerful laptop, so now I'm using the antique one :(

OpenStudy (mertsj):

It's C

OpenStudy (anonymous):

merts are you sure. when i tried this i came up with b

OpenStudy (anonymous):

can you demonstrate please?

OpenStudy (anonymous):

Sorry to hear about that Cholorophyll. I broke my laptop as well so I have to walk on campus to use the schools computer for everything

OpenStudy (anonymous):

"walk" , for real in this freezing weather... :'(

OpenStudy (mertsj):

Bottom times derivative of top is \[x ^{\frac{1}{2}}(2x-3-\frac{1}{2}x ^{\frac{-1}{2})}\]

OpenStudy (anonymous):

I live like a 5 mins distance plus my roomate somtimes give me a ride. not so bad

OpenStudy (mertsj):

-top times derivative of bottom IS: \[-(x^2-3x-x ^{\frac{1}{2}})(\frac{1}{2}x ^{\frac{-1}{2}})\]

OpenStudy (mertsj):

Put those two together and you have your numerator.

OpenStudy (mertsj):

Put all of that over x which is the denominator squared

OpenStudy (mertsj):

Now let's distribute and simplify

OpenStudy (anonymous):

That's correct, so you take denominator is 2√x from inside the parentheses, then distribute with √x out side

OpenStudy (anonymous):

I don't see any flaw in my calculation!

OpenStudy (mertsj):

\[2x ^{\frac{3}{2}}-3x ^{\frac{1}{2}}-\frac{1}{2}x ^{0}-\frac{1}{2}x ^{\frac{3}{2}}+\frac{3}{2}x ^{\frac{1}{2}}+\frac{1}{2}x ^{0}\]

OpenStudy (mertsj):

remember this is all over x

OpenStudy (mertsj):

Now collect the like terms.

OpenStudy (mertsj):

\[\frac{3}{2}x ^{\frac{3}{2}}-\frac{3}{2}x ^{\frac{1}{2}}\]

OpenStudy (mertsj):

Remember that is all over x so now divide each term by x

OpenStudy (mertsj):

\[\frac{3}{2}x ^{\frac{1}{2}}-\frac{3}{2}x ^{\frac{-!}{2}}\]

OpenStudy (mertsj):

Choice C

OpenStudy (anonymous):

ok thanks

OpenStudy (mertsj):

Ok. Now I'll check it with Wolf

OpenStudy (anonymous):

What the result, Mertsj?

OpenStudy (anonymous):

Is it 3 (x -1)/ 2 √x

OpenStudy (mertsj):

\[\frac{3\sqrt{x}}{2}-\frac{3}{2\sqrt{x}}\]

OpenStudy (mertsj):

Which is exactly what I said albeit in a slightly different form.

OpenStudy (mertsj):

Your answer is the same thing chlorophyll

OpenStudy (anonymous):

Yep, just in add together, --> (3x -1)/ 2√x

OpenStudy (mertsj):

That would be 3x-3

OpenStudy (anonymous):

I mean 3(x -1)/ 2√x

OpenStudy (anonymous):

Happy, Jinnie :)